OpenStudy (anonymous):

Find all critical numbers for the function f(x)=9x3−27x+4, and then list them (separated by commas) in the box below. If there are no critical numbers, enter None.

OpenStudy (anonymous):

@Chlorophyll can you give me a hand?

OpenStudy (anonymous):

Critical numbers is obtained when let y' = 0 and y" = 0

OpenStudy (anonymous):

all of them, so all i need to do is find the zeros of the derivative?

OpenStudy (anonymous):

so for this question 27 (x^2-1) x=1, and x=-1

OpenStudy (anonymous):

so then i plug that into original equation and find y?

OpenStudy (anonymous):

so 9(1)^3−27(1)+4=-14 and 9(-1)^3−27(-1)+4=22

OpenStudy (anonymous):

so (-1,22) and (1,-14) are critical numbers?

OpenStudy (anonymous):

how would i find y"?

OpenStudy (anonymous):

so again roots of y; are 1 and -1

OpenStudy (anonymous):

and root of y'' =0

OpenStudy (anonymous):

idk doesnt sound familiar so im guessing not

OpenStudy (anonymous):

so what is my answer?

OpenStudy (anonymous):

it says that my answer isnt a number its a list of numbers

OpenStudy (anonymous):

so try that as an answer?

OpenStudy (anonymous):

because that didnt work either

OpenStudy (anonymous):

idk just telling you what i was told

OpenStudy (anonymous):

Maybe you should link me to the source page or your post is incorrect!

OpenStudy (anonymous):

brb got to go eat

OpenStudy (anonymous):

This is not the same problem! ( Yes, late dinner time here, too! )

OpenStudy (anonymous):

Try x = 3

OpenStudy (anonymous):

i figured it out it was 1,-1

OpenStudy (anonymous):

Oh, they means critical NUMBERS are just X values, (without y values ), Hmm!

OpenStudy (anonymous):

ya i guess so

OpenStudy (anonymous):

yup but that first link that i sent you i cant get it for some reason i got 3 as an answer

OpenStudy (anonymous):

You had the answer x= 3?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

x=-3 sorry

OpenStudy (anonymous):

Mine is x = 3

OpenStudy (anonymous):

didnt work either

OpenStudy (anonymous):

Yes, your x = -3 is correct!

OpenStudy (anonymous):

i guess it was none for some reason